Abin contains six parts. Two of the parts are defective and four are acceptable. If three of the six parts are selected from the bin, how large is the sample space? which counting rule did you use and why? For this sample apace, what is the probability that exactly one of the three sampled parts is defective?
Answer by stanbon(75887)

Abin contains six parts. Two of the parts are defective and four are acceptable.
If three of the six parts are selected from the bin, how large is the sample space?
0 defective:: 4C3 = 4 unique sets of 3 acceptable
1 defective: 2C1*4C2 = 2*6 = 12 unique sets of 1 defective and 2 acceptable
2 defective: 2C2*4C1 = 1*4 = 4 uniqe sets of 2 defective ad 1 acceptable
-----------------
which counting rule did you use and why?
Multiplication Rule
-------
For this sample apace, what is the probability that exactly one of the three sampled parts is defective?
Ans: 12/20 = 3/5
======================
Cheers,
Sta H.